OpenStudy (anonymous):

secB/sinB - sinB/cosC = cotB if someone could solve i would really appreciate it please :)

OpenStudy (anonymous):

Rewrite the sec and cot as their equivalent sin and cosine formula

OpenStudy (anonymous):

use sine and cosine

OpenStudy (anonymous):

sec= 1/cos cot=1/tan tan=sin/cos Use that information to simplify it further

OpenStudy (anonymous):

(1/cosB)/(sinB) - sinB/cosB = cotB

OpenStudy (anonymous):

cotB=1/tanB ??

OpenStudy (anonymous):

I thought the original problem had sin B/cos C, which you can't combine together in any way

OpenStudy (anonymous):

its cosB

OpenStudy (anonymous):

So yes it's true. You are correct